psirockin2:

Help with a math problem? It's Matrices and stuff (Seriously please help... MyMathLab is mean)

psirockin2:

1 attachment
psirockin2:

That's the question

psirockin2:

I know that x+y+z=40,000 and that 0.02x+0.04y+0.06z=1,520 and that z=x/2 I just don't know what to do with them

psirockin2:

Do I turn each of those into a matrix and type them into the calculator?

imqwerty:

you can solve system of linear equations with matrices